OpenStudy (anonymous):

Medal + Fan :) The value of x + x(^x) when x = 2 is: (a) 10, (b) 16, (c) 18, (d) 36, (e) 64 btw, I completely regret my name :)

OpenStudy (anonymous):

[x + x(x ^{x})\]) put the value of x = 2 in the above expression we get 2 + 2(22) = 2 + 2(2 × 2) = 2 + 2(4) = 2 + 8 = 10 so your answer is a :) hope that helped you!

OpenStudy (anonymous):

ah, woops I messed up on the equation part, haha xD

OpenStudy (anonymous):

thank you!!